PLEASE ANSWER NUMBER 54. (1 point) Evaluate the triple integral on the given domain slf (x² + y2 +22)3/2 dxdydz where G={(x,y,z): x² + y2 +z? <4} 5
Answer #1

5. Z=16-a2 ye z=a²+ ye 2²ty&= 16-2²-ye 2a²+&ge=16 a&ty? =8=cereje Polao coordinates =300gb , y=osino de&xg8 secosa toß sin?a=-Qo3Jdodo - jere. 0116_202) dode par lest [166203 Jetode tjera a de do dododno of Chinese– 378133 mg = [ 818) – bu ]de - 32 (